OpenStudy (anonymous):

HELP! Simplify 6√10+9√10

OpenStudy (anonymous):

82.704

jagr2713 (jagr2713):

@williambaltus please avoid giving direct answers http://openstudy.com/code-of-conduct This site has a policy for "Give Help, Not Answers" violations of such policies could potentially result in an account suspension or ban.

jagr2713 (jagr2713):

\[Add 6\sqrt{10} and 9\sqrt{10} \to get 15\sqrt{10}\]

OpenStudy (anonymous):

Thank you so much for the help. It means a lot. I appreciate it!

jagr2713 (jagr2713):

Anytime

OpenStudy (anonymous):

Oh my bad Brooke, for this one you simply sqaure root 19 and than sqaure root 10, after that you mulitply all answers and you get it that is all.